The equation of Circle P is X squared + Y squared - 6x + 4y= 21, and the equation of circle Q is X squared + y squared - 8x + 10y = 11. How do I find the difference between the centers of these circles?
Answer by scott8148(6628)

for a circle of the form x^2+ax+y^2+by+c=d; the coordinates of the center are [(-a/2),(-b/2)]

circle P is centered at (3,-2) and circle Q is centered at (4,-5)

if by difference, you mean distance; then d=sqrt[(4-3)^2+(-5-(-2))^2] __ d=sqrt(10)